A. A researcher reports that the average salary of College Deans is more than $63,000. At a = 0.01 level of confidence, test the claim that the College Deans earn more than $63,000 a month.

1
Expert's answer
2022-06-15T13:03:41-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le63000"

"H_1:\mu>63000"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a right-tailed test is "z_c = 2.3263."

The rejection region for this right-tailed test is "R = \{z:z>2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{65700-63000}{5250/\sqrt{35}}=3.0426"

Since it is observed that "z=3.0426>2.3263=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(z>3.0426)=0.001173," and since "p=0.001173<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 63000, at the "\alpha = 0.01" significance level.